OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2 y – 9 = 2(x – 5) y – 5 = 2(x + 9) y + 9 = 2(x – 5) y – 9 = 2(x + 5)

OpenStudy (anonymous):

\[(y-y_1)=m(x-x_1)\]

OpenStudy (anonymous):

Point slope form is \[\large y-y_{1}=m(x-x_{1})\] with m as the slope, and \(x_{1}, y_{1}\) being the coordinates

OpenStudy (anonymous):

Great

OpenStudy (nayef):

The point: (a,b) ,The slpe: m , so the equation of the line pass through the point and has the slope m is (y-b)=m(x-a).